06/04/99 owl shift summary

Angela Bellavance (bellavan@rice.edu)
Fri, 4 Jun 1999 07:11:34 -0500 (CDT)

Shift Summary for owl shift, 06/04/99

On Shift: Bob T.
Angela B.

Hours of beam delivered at all intensities: 4
Hours of beam delivered at nominal intensity (1.2E12): 4
Hours of E799 data: 3
Hours of special runs: 0

Accounting of lost time: DA problems

===================================================================
Runs ended this shift:

13342 - BEAM_799 - 1.2 M
13343 - BEAM_799 - not usable
13344 - BEAM_799 - 4.0 M
13345 - BEAM_799 - in progress

Tapes brought to FCC: 0

===================================================================
Accesses during shift: None

===================================================================
Description of shift:

run time comment
-----------------------
13342 00:00 Inherit run from evening shift, getting beam.
" 00:17 Begin experiencing stream 5 problems
" 00:22 Stop run. Reset laser repeater (power cycle).
13343 00:42 Notice that SDAQ HV monitoring is frozen. A restart of
that process doesn't work. Will wait for loss of beam to
try an SDAQ fresh start. SteveB. and Emmanuel are
struggling with the DA. This is a test run with no useful
info.
13344 ~01:20 Start new run with 150MeV magnet field.
" 01:52 Message: Tuning MCenter - no beam to us.
" 02:11 Checking online plots. Number of minor problems -
individuals are emailed about them (RickK, JasonL,
Elizabeth, Val, Julie, SashaL, Leo, JasonH, Sydney,
SteveB, RickT, Valeri).
" 03:21 Beam returns.
" 04:08 Message: Beam to users (MCenter tuning over)
" 04:33 Stopping run to go to new magnet setting.
13345 04:45 Start new run with magnet at 100MeV.
" 05:18 Beam go bye-bye.
" 05:31 Message: TeV quench. Recovery underway.
" 05:47 Did an SDAQ fresh start to try to get HV monitoring window
working again. Window restarts, but immediately freezes
again. Theo emailed.
" 06:27 Message: TeV QPM problems. Estimate 45 min.
" 07:00 Hand off to day shift.

General problems that need to be looked into:
- FPED process is running too slow and too frequently causing every 4th or
5th spill to be dead.
- Stream 5 is sick. Laser repeater has to be power cycled every so often.
- Laser and ped triggers are occuring _in_ spill at a low level -> causes
significant monitoring problems.
- We are out of printer paper. Header pages were dug out of recycling and
we are using the backside of them for now. Get more from the high-rise?

Regards, Angela & Bob T.

_______
Angela Bellavance
bellavan@rice.edu home: (630) 548-3412
http://www.owlnet.rice.edu/~bellavan/ work: (630) 840-4144
"It has yet to be proven that intelligence has any survival value."
- Arthur C. Clarke